can you tell me how to add and subtract fractions? here is an easy example I dont quite understand. i couldnt find one to subtract in order to practice.
2/3 + 3/4=
Thank You.
Liz

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

convert them to the same denominator
2/3 +3/4
multiply 3*4=12
the first fraction is 2/3
12/3=4
so multiply both numeration and denominator by 4 that is by 4/4 which is really 1 in disguise
4/4*2/3=8/12
next fraction
12/4=3
multiply 3/4*3/3=9/12
so now we have 8/12+9/12
add 8+9=17
17/12
now reduce the improper fraction to a mixed number
17/12= 1 5/12 done
so we did
2/3+3/4
3*4=12
2/3*4/4+3/4*3/3
8/12+9/12
17/12
1 5/12

Answer by Alan3354(69443)

2/3 + 3/4
------------
Step 1 is to get the same denominator in all fractions.
One way that always works is to multiply the the DENs given:
3*4 = 12
Use 12
Then change each fraction by multiplying NUM and DEN to get 12 in the DEN:
2/3 = 8/12
3/4 = 9/12
---------
Then you can add them
8/12 + 9/12 = (8+9)/12
= 17/12
-------
Sometimes you're asked to reduce it
--> 1 5/12
--------------